# Edge densities of drawings of graphs with one forbidden cell

### v2

@@ 6 Quasiplanar and [Uncaptioned image]-Free Drawings
For quasiplanar drawings, we improve the lower bound for non-homotopic drawings of simple graphs from 7n287n-28 [kaufmann_density_formula] to 7.5n287.5n-28 (Construction˜7). We remark that there still remains a gap to the upper bound of 8n208n-20 [ackerman_quasiplanar].
For [Uncaptioned image]-free drawings, we improve the lower bound for non-homotopic drawings of simple graphs from $7.5n-\mathcal{O}(1)$ [ackerman_quasiplanar] to 8n288n-28 (Construction˜8), almost matching the upper bound of 8n208n-20 [ackerman_quasiplanar]. For [Uncaptioned image]-free simple drawings of graphs, we give a lower bound of 7n307n-30 (Construction˜6) and prove a near-tight upper bound of 7n207n-20 (Theorem˜6.1). The previous upper bound was 8n208n-20 [ackerman_quasiplanar] in this case, too. See again Table˜2.
@@
quasiplanar     multigraphNon-homotopic         [kaufmann_density_formula]8n208n-20       [ackerman_quasiplanar]8n208n-20
    simple graphNon-homotopic           implicitly in [kaufmann_density_formula]7n287n-28           Construction˜77.5n287.5n-28       [ackerman_quasiplanar]8n208n-20
Simple drawing         [kaufmann_density_formula]6.5n206.5n-20       [ackerman_quasiplanar]6.5n206.5n-20
[Uncaptioned image]-free     multigraphNon-homotopic         [kaufmann_density_formula]8n208n-20       [ackerman_quasiplanar]8n208n-20
@@
Simple drawing                     Construction˜6 based on [ackerman_quasiplanar]7n307n-30          Theorem˜6.17n207n-20       [ackerman_quasiplanar]8n208n-20
@@
6.1 Upper Bound for [Uncaptioned image]-Free Drawings of Simple Graphs
Ackerman and Tardos proved in [ackerman_quasiplanar] that any $n$-vertex quasiplanar drawing ($n\geq 3$) contains at most 8n208n-20 edges, but explicitly mention that they only use [Uncaptioned image]-freeness. For arbitrary (not necessarily non-homotopic) drawings, this would contradict Construction˜3. However, their proof relies on the additional assumption that the drawing is crossing-minimal quasiplanar, which they use to rule out cells $c$ with $\left\lVert c\right\rVert\leq 2$. In arbitrary crossing-minimal [Uncaptioned image]-free drawings, this cannot be concluded [^4] Consider e.g. a [Uncaptioned image]-cell $c$ and add an edge $e$ that enters and leaves $c$ through the same edge $f$ on the boundary of $c$. Then, trying to remove the [Uncaptioned image]-cell formed by $e$ and $f$ by redrawing reintroduces $c$.. But recall from Section˜2 that in any non-homotopic drawing every cell has size at least $3$. Therefore, the proof of the 8n208n-20 upper bound in [ackerman_quasiplanar] applies to non-homotopic [Uncaptioned image]-free drawings, even of multigraphs.
Ackerman and Tardos also considered the case of simple quasiplanar drawings, for which they obtained an upper bound of 6.5n206.5n-20 edges via a similar proof idea. The final case distinction in this proof however, works only for quasiplanar drawings and not for [Uncaptioned image]-free drawings. Here, we extend their proof to simple [Uncaptioned image]-free drawings by adding one further case. This yields the slightly higher upper bound of 7n207n-20. In fact, our bound is best possible (up to an additive constant), as we shall later construct simple $n$-vertex graphs with [Uncaptioned image]-free drawings and 7n307n-30 edges (Construction˜6). In particular, the edge densities of simple quasiplanar drawings and simple [Uncaptioned image]-free drawings are indeed different.
As mentioned above, we modify the proof of Ackerman and Tardos. Hence, we only sketch this proof and refer to [ackerman_quasiplanar] for more details.
Theorem 6.1.
Let $\Gamma$ be a [Uncaptioned image]-free drawings of an $n$-vertex multigraph $G$. If $\Gamma$ is non-homotopic and $n\geq 3$, then $G$ has at most 8n208n-20 edges. If $\Gamma$ is simple and $n\geq 4$, then $G$ has at most 7n207n-20 edges.
Proof sketch: Consider any [Uncaptioned image]-free drawing $\Gamma$ of an $n$-vertex multigraph. We later distinguish whether $\Gamma$ is non-homotopic or simple.
Initially, give each cell charge $\text{ch}(c)\coloneqq\left\lVert c\right\rVert-4$. By Lemma˜2.1, the total charge is therefore 4n84n-8. We will redistribute the charges in two steps and, via double counting, obtain an upper bound on the number of edges. In the first discharging step, we obtain new charges ch by setting $\text{ch${}^{\prime}$}(c)\coloneqq\text{ch}(c)+\frac{1}{5}$ for each [Uncaptioned image]-cell $c$ and $\text{ch${}^{\prime}$}(c)\coloneqq\text{ch}(c)-\frac{e_{\rm in}(c)}{5}$ for each $c\in\mathcal{C}_{\geq 5}$. By Lemma˜2.2, the total charge does not increase in this step. Using [Uncaptioned image]-freeness, it can be checked via a case distinction that for every cell $c$, we have $\text{ch${}^{\prime}$}(c)\geq\frac{v(c)}{5}$.
The final charge distribution ch′′ is obtained by moving charge from cells to vertices. Namely, we distribute the excess charge $\text{ch${}^{\prime}$}(c)-\frac{v(c)}{5}$ of each cell fairly among its $v(c)$ incident vertices. We lower bound the charge $\text{ch${}^{\prime\prime}$}(v)$ of an arbitrary vertex $v$ by a case distinction on the type of the cell $c_{v}$ that is obtained when removing $v$ and its incident edge segments (see Figure˜5, left).
Figure 5: Left: Removing a vertex $v$ and its incident edge segments, a new cell $c_{v}$ (shaded gray) is formed. Others: Case distinction in Theorem˜6.1 where $c_{v}$ is a Refer to caption-cell.
In the non-homotopic setting for $n\geq 3$, Ackerman and Tardos prove $\text{ch${}^{\prime\prime}$}(v)\geq\frac{4}{5}$. The minimum is obtained if $c_{v}$ is a [Uncaptioned image]-cell. When $\Gamma$ is simple quasiplanar and $n\geq 4$, they show $\text{ch${}^{\prime\prime}$}(v)\geq\frac{7}{5}$ and equality can be obtained when $c_{v}$ is a [Uncaptioned image]-cell. To complete the case distinction in our setting of simple [Uncaptioned image]-free $\Gamma$, it suffices to consider the additional case that $c_{v}$ is a [Uncaptioned image]-cell: Note that $\text{ch${}^{\prime\prime}$}(v)$ is unchanged, whether an edge segment of $c_{v}$ is crossed once or more than once, since additional crossings only increase the number of [Uncaptioned image]-cells incident to $v$, which cannot send any charge to $v$. Thus, we only have to compute $\text{ch${}^{\prime\prime}$}(v)$ if $s=1,2,3$ edge segments of $c_{v}$ are crossed (for $s=0$, $\Gamma$ is disconnected), see Figure˜5, second-to-left to right. If $s=1$, $v$ has final charge $\frac{10}{5}$, sent from a cell of size $7$ with one incident vertex. If $s=2$, $\text{ch${}^{\prime\prime}$}(v)=\frac{6}{5}+\frac{2}{5}$ given by a cell of size 6 with one incident vertex and a [Uncaptioned image]-cell. Finally, if $s=3$, $\text{ch${}^{\prime\prime}$}(v)=3\cdot\frac{2}{5}$ from the three incident [Uncaptioned image]-cells. Thus in total, $\text{ch${}^{\prime\prime}$}(v)\geq\frac{6}{5}$ for every vertex $v$. Finally, by rearranging
$4n-8=\sum_{c\in C}\text{ch}(c)\geq\sum_{c\in C}\frac{v(c)}{5}+\sum_{v\in V}\text{ch${}^{\prime\prime}$}(v)\geq\frac{2}{5}|E|+|V|\cdot\min_{v}\text{ch${}^{\prime\prime}$}(v),$
we obtain $|E|\leq 8n-20$ for non-homotopic and $|E|\leq 7n-20$ for simple [Uncaptioned image]-free drawings. ∎
@@ 6.2 Lower Bounds for [Uncaptioned image]-Free and Quasiplanar Drawings of Simple Graphs
Let us start with simple [Uncaptioned image]-free drawings. We give a lower bound construction that is tight up to an additive constant. In fact, in this case, it is enough to take a subdrawing of the construction in [ackerman_quasiplanar, Theorem 4].
@@
For infinitely many values of $n$, there is a simple $n$-vertex graph with 7n307n-30 edges that admits a [Uncaptioned image]-free drawing.
Proof 6.2.
Fix $m\geq 1$ and start with an $m\times 3$-grid of hexagons on $n=8m+6$ vertices, each hexagon containing a straight-line drawing of $K_{6}\setminus e$ where the vertical diagonal is missing (see Figure˜6, left). Now, identify opposite vertices along the side of the grid of length $m$ to obtain the surface of a cylinder whose top and bottom boundary contain 6 vertices. In the resulting drawing, every vertex not on the boundary has degree 11. Now, add long edges as depicted in the center of Figure˜6 whenever this is possible. This increases the degree of every vertex sufficiently far away from the boundary to 14. After adding three uncrossed edges each on the top and bottom (see Figure˜6, right), we arrive at the final drawing $\Gamma$. Counting vertex degrees near the boundary, it can be confirmed that the number of edges in $\Gamma$ is 7n307n-30.
Figure 6: Illustrations of Construction˜6.
Before adding the long edges, the drawing is clearly [Uncaptioned image]-free and no [Uncaptioned image]-cells are created by adding the long edges and the six uncrossed edges. Finally, the simplicity of $\Gamma$ follows from the fact that $\Gamma$ is locally homeomorphic to a straight-line drawing and no two edges are long enough to meet twice on the cylinder.
In the remainder of this section, we construct non-homotopic drawings of simple graphs. We start with a quasiplanar construction, followed by a construction of [Uncaptioned image]-free such drawings.
@@
Proof 6.3.
Let $n\geq 14$ be even. We describe a drawing on $n$ vertices with the desired properties. Precise edge paths are obtained by avoiding unnecessary crossings. Figure˜7 (Top) depicts the final drawing $\Gamma$. We draw
  1. 1.
    an $\frac{n}{2}$-cycle $C_{\rm in}$ as a circle with equidistant points.
  2. 2.
    a concentric copy $C_{\rm out}$ of $C_{\rm in}$ with larger diameter.
  1. 3.
    a straight edge between any pair of partners.
  2. 4.
    for any vertex $v$, a straight edge to $v^{\prime}_{+1}$.
  3. 5.
    for any vertex $v$ on $C_{\rm in}$ an edge from $v$ to $v_{+2}$, drawn through the interior of $C_{\rm in}$.
  4. 6.
    for any vertex $v$ on $C_{\rm out}$ an edge from $v$ to $v_{+2}$, drawn through the exterior of $C_{\rm out}$.
  5. 7.
    for any vertex $v$ on $C_{\rm in}$, an edge to $v^{\prime}_{+2}$. The edge passes through the space between $C_{\rm in}$ and $C_{\rm out}$, crosses $C_{\rm out}$ in the edge $v^{\prime}v^{\prime}_{+1}$ and continues through the exterior of $C_{\rm out}$.
  6. 8.
    for any vertex $v$ on $C_{\rm out}$, an edge to $v^{\prime}_{+2}$. Edges are drawn symmetrically to the last step.
  7. 9.
    for any vertex $v$ on $C_{\rm in}$ an edge to $v_{+3}$. The edge first passes through the space between $C_{\rm in}$ and $C_{\rm out}$, crosses $C_{\rm in}$ between $v_{+1}$ and $v_{+2}$ and then goes through the interior of $C_{\rm in}$.
  8. 10.
    an edge between $v$ and $v_{+3}$ for any vertex $v$ on $C_{\rm out}$, symmetrically to the last step.
  9. 11.
    a zig-zag path inside of $C_{\rm in}$ whose edges connect vertices of distance at least 4 on $C_{\rm in}$.
  10. 12.
    another different zig-zag path in the interior of $C_{\rm in}$ such that no parallel edges are created.
  11. 13.
    a zig-zag path outside of $C_{\rm out}$, again connecting vertices of distance at least 4 on $C_{\rm out}$.
  12. 14.
    another such zig-zag path in the exterior of $C_{\rm out}$.
Figure 7: The drawings of Constructions˜7 and 8 with the outer zig-zag paths omitted for visual clarity, as well as zoomed-in strips of the local edge patterns.
Top: Construction˜7 where edges except on the inner zig-zag paths are partitioned by color into three plane drawings. The edges that form lenses with the thick edge are dashed.
Bottom: Construction˜8 where green edges are those not present in Construction˜7. The shaded area is bounded by three pairwise crossing edges.
The first three steps add $\frac{n}{2}$ edges each, step˜4 adds $n$ edges, the six steps thereafter $\frac{n}{2}$ edges each, and the last four steps $\frac{n}{2}-7$ edges each, totaling the desired 7.5n287.5n-28 edges. As the various steps connect vertices at different distances, the underlying graph is simple.
We now argue that $\Gamma$ is quasiplanar. First, consider all edges except those in zig-zag paths. These can be locally 3-colored such that one color class consists only of the edges in step˜3 and edges of the same color do not cross (see Figure˜7 ). Therefore, any 3 pairwise crossing edges would contain an edge from step˜3, but each such edge is only crossed by two other edges, which in turn do not cross.
Now, consider without loss of generality the zig-zag paths $P$ and $P^{\prime}$ in the interior of $C_{\rm in}$ and some edge $e=uv$ in $P$. Observe that $e$ is crossed by three sets of edges: (i) Edges in $P^{\prime}$, (ii) edges from steps 1-10 that cross $e$ at $u$, and (iii) edges from steps 1-10 that cross $e$ at $v$.
Finally, we argue that $\Gamma$ is non-homotopic. Note that in the above iterative construction, simplicity is only violated by the edges added in steps˜9 and 10. Each such edge is part of four lenses (see dashed edges in Figure˜7 (Top)), all of which are non-empty.
@@
Proof 6.4.
Let $n\geq 14$ be even and start with the [Uncaptioned image]-free drawing from Construction˜7. We will reuse the notation established there. For any vertex $v$ on $C_{\rm out}$, we draw an edge from $v$ to $v^{\prime}_{+3}$ that first moves through the space between $C_{\rm in}$ and $C_{\rm out}$, crosses $C_{\rm in}$ in $v^{\prime}_{+1}v^{\prime}_{+2}$ and then passes through the interior of $C_{\rm in}$. The resulting drawing is shown in Figure˜7 (Bottom).
Checking along the paths of the newly added edges, it can be verified that the drawing remains [Uncaptioned image]-free. Note that this process adds $\frac{n}{2}$ edges, none of which are parallel to each other or previous edges. Therefore, we obtain the claimed density.
6.3 Non-Quasiplanar [Uncaptioned image]-Free Simple Drawings

### v3

@@ 5 Quasiplanar and [Uncaptioned image]-Free Drawings
quasiplanar     multigraphNon-homotopic            [kaufmann_density_formula, Thm. 6.7]8n208n-20           [ackerman_quasiplanar, Thm. 3]8n208n-20
    simple graphNon-homotopic              [kaufmann_density_formula, Thm. 6.7] (implicit)7n287n-28           Construction˜67.5n287.5n-28           [ackerman_quasiplanar, Thm. 1]8n208n-20
Simple drawing            [kaufmann_density_formula, Thm. 6.11]6.5n206.5n-20           [ackerman_quasiplanar, Thm. 5]6.5n206.5n-20
[Uncaptioned image]-free     multigraphNon-homotopic            [kaufmann_density_formula, Thm. 6.7]8n208n-20           [ackerman_quasiplanar, Thm. 3]8n208n-20
@@
Simple drawing                       Construction˜8 based on [ackerman_quasiplanar, Thm. 4]7n307n-30           [ackerman_quasiplanar, Thm. 6]7n207n-20
@@
For quasiplanar drawings, we improve the lower bound for non-homotopic drawings of simple graphs from 7n287n-28 [kaufmann_density_formula] to 7.5n287.5n-28 (Construction˜6). We remark that there still remains a gap to the upper bound of 8n208n-20 [ackerman_quasiplanar].
@@
For [Uncaptioned image]-free drawings, we improve the lower bound for non-homotopic drawings of simple graphs from $7.5n-\mathcal{O}(1)$ [ackerman_quasiplanar] to 8n288n-28 (Construction˜7), almost matching the upper bound of 8n208n-20 [ackerman_quasiplanar].
For [Uncaptioned image]-free simple drawings, we give a lower bound of 7n307n-30 (Construction˜8) based on a construction of Ackerman and Tardos , who also proved an upper bound of 7n207n-20 [ackerman_quasiplanar, Theorem 6]. See again Table˜2.
@@ 5.1 Lower Bounds for [Uncaptioned image]-Free and Quasiplanar Drawings of Simple Graphs
We start with two constructions for non-homotopic drawings of simple graphs. First, a quasiplanar construction, followed by a construction of [Uncaptioned image]-free drawings.
@@
Proof 5.1.
Let $n\geq 14$ be even. We describe a drawing on $n$ vertices with the desired properties. Precise edge paths are obtained by avoiding unnecessary crossings. Figure˜5 (top) depicts the final drawing $\Gamma$. We draw
  1. 1.
    an $\frac{n}{2}$-cycle $C_{\rm in}$ as a circle with equidistant points and a concentric copy $C_{\rm out}$ of $C_{\rm in}$ with larger diameter.
  1. 2.
    a straight edge between any pair of partners,
  2. 3.
    for any vertex $v$, a straight edge to $v^{\prime}_{+1}$,
  3. 4.
    for any vertex $v$ on $C_{\rm in}$ ($C_{\rm out}$) an edge from $v$ to $v_{+2}$, drawn through the interior of $C_{\rm in}$ (exterior of $C_{\rm out}$),
  4. 5.
    for any vertex $v$ on $C_{\rm in}$ ($C_{\rm out}$), an edge to $v^{\prime}_{+2}$. The edge passes through the space between $C_{\rm in}$ and $C_{\rm out}$, crosses $C_{\rm out}$ ($C_{\rm in}$) in the edge $v^{\prime}v^{\prime}_{+1}$ and continues through the exterior of $C_{\rm out}$ (interior of $C_{\rm in}$),
  5. 6.
    for any vertex $v$ on $C_{\rm in}$ ($C_{\rm out}$) an edge to $v_{+3}$. The edge first passes through the space between $C_{\rm in}$ and $C_{\rm out}$, crosses $C_{\rm in}$ ($C_{\rm out}$) between $v_{+1}$ and $v_{+2}$ and then goes through the interior of $C_{\rm in}$ (exterior of $C_{\rm out}$),
  6. 7.
    two zig-zag paths inside of $C_{\rm in}$ (outside of $C_{\rm out}$), whose edges connect vertices of distance at least 4 on $C_{\rm in}$ ($C_{\rm out}$) such that no parallel edges are created.
Figure 5: The drawings of Constructions˜6 and 7 for $n=24$ with the outer zig-zag paths omitted for visual clarity, as well as zoomed-in strips of the local edge patterns.
Top: Construction˜6 where edges except on the inner zig-zag paths are partitioned by color into three plane drawings. In the zoomed-in strip, the incident edges of one vertex are marked by the step of the construction, in which they are added. In the The edges that form lenses with the thick edge are dashed.
Bottom: Construction˜7 where green edges are those not present in Construction˜6. The shaded area is bounded by three pairwise crossing edges (drawn dashed).
Step 1 adds $n$ edges, Step 2 adds $\frac{n}{2}$ edges, Steps 3-6 add $n$ edges each, and the final step contributes another $4\cdot(\frac{n}{2}-7)$ edges, totaling the desired 7.5n287.5n-28 edges. As the various steps connect vertices at different distances, the underlying graph is simple.
We now argue that $\Gamma$ is quasiplanar. First, consider all edges except those in zig-zag paths. These can be locally 3-colored such that one color class consists only of the edges in step˜2 and edges of the same color do not cross (see Figure˜5). Therefore, any 3 pairwise crossing edges would contain an edge from step˜2, but each such edge is only crossed by two other edges, which in turn do not cross.
Now, consider without loss of generality the zig-zag paths $P$ and $P^{\prime}$ in the interior of $C_{\rm in}$ and some edge $e=uv$ in $P$. Observe that $e$ is crossed by three sets of edges: (i) Edges in $P^{\prime}$, (ii) edges from Steps 1-10 that cross $e$ near $u$, and (iii) edges from Steps 1-10 that cross $e$ near $v$.
Finally, we argue that $\Gamma$ is non-homotopic. Note that in the above step-by-step construction, simplicity is only violated by the edges added in step˜6. Each such edge is part of four lenses (see dashed edges in Figure˜5 (top)). As each of these lenses contains a vertex, the drawing is non-homotopic.
@@
Proof 5.2.
Let $n\geq 14$ be even and start with the non-homotopic quasiplanar and hence also [Uncaptioned image]-free drawing from Construction˜6. We will reuse the notation established there. For any vertex $v$ on $C_{\rm out}$, we draw an edge from $v$ to $v^{\prime}_{+3}$ that first moves through the space between $C_{\rm in}$ and $C_{\rm out}$, crosses $C_{\rm in}$ in $v^{\prime}_{+1}v^{\prime}_{+2}$, and then passes through the interior of $C_{\rm in}$. This process adds $\frac{n}{2}$ edges, none of which are parallel to each other or previous edges. Therefore, we obtain a drawing of a simple graph with the claimed density. The resulting $\Gamma$ is shown in Figure˜5 (Bottom).
Note that $\Gamma$ is not quasiplanar as it contains triples of pairwise crossing edges (see again Figure˜5 (Bottom)). However, checking the cells along the paths of the newly added edges, it can be verified that the construction preserves [Uncaptioned image]-freeness. Similarly, one can check that none of the lenses introduced by the new edges is empty, and so $\Gamma$ also remains non-homotopic.
Finally, we consider simple [Uncaptioned image]-free drawings. In this setting, Ackerman and Tardos mention that a $7n-\mathcal{O}(1)$ lower bound can be obtained by taking a subdrawing of their construction in [ackerman_quasiplanar, Theorem 4]. However, they neither specify nor optimize the constant term. Doing so, we obtain a lower bound of 7n307n-30, almost matching the 7n207n-20 upper bound by Ackerman and Tardos.
@@
For infinitely many values of $n$, there is a simple $n$-vertex graph with 7n307n-30 edges that admits a simple [Uncaptioned image]-free drawing.
Proof 5.3.
Fix $m\geq 3$ and start with an $m\times 3$-grid of hexagons on $8m+6$ vertices, each hexagon containing a straight-line drawing of $K_{6}\setminus e$ where the vertical diagonal is missing (see Figure˜6, left). Identify opposite vertices along the side of the grid of length $m$ to obtain the surface of a cylinder whose top and bottom boundary contain six vertices each. This leaves us with $n\coloneqq 6m+6$ vertices, which we group into $2m+2$ layers of three vertices each, labeled from top to bottom, see Figure˜6, middle. The vertices on the boundary are those in layers $1$ and $2$ (top boundary), and $2m+1$ and $2m+2$ (bottom boundary). All vertices in layers $3,\ldots,2m$ currently have degree $11$, while the vertices in layers $2$ and $2m+1$ have degree $9$ and those in layers $1$ and $2m+2$ degree $4$. Note that, so far, the drawing is [Uncaptioned image]-free.
Now, add long edges as depicted in Figure˜6, middle. Every vertex gains one edge to a neighboring layer and one edge each to a layer three above and below (if those layers exist).
This increases the degree of the vertices on layers $4,\ldots,2m-1$ by $3$ and the remaining vertex degrees by $2$. After adding a triangle on layer $1$ and $2m+2$, we arrive at the final drawing $\Gamma$ (see Figure˜6, right). Grouping layers with the same vertex degrees, we find that the number of edges in $\Gamma$ is
$\frac{1}{2}\Big((2m-4)\cdot 3\cdot 14+2\cdot 3\cdot(13+11+7)\Big)=42m+12=7n-30.$
Figure 6: Illustrations of Construction˜8 for $m=4$.
Note that no [Uncaptioned image]-cells are created by adding the long edges and the two triangles. Finally, the simplicity of $\Gamma$ follows from the fact that $\Gamma$ is locally homeomorphic to a straight-line drawing and no two edges are long enough to meet twice on the cylinder.
5.2 Simple Non-Quasiplanar [Uncaptioned image]-Free Drawings